STATISTICS: What is the probability of getting 3 heads on 12 tosses of an un fair coin p(H)= 0.7?
Answer #1

P(Head)=0.7

Probability of getting 3 heads= P(H1)*P(H2)*P(H3)=0.7*0.7*0.7=0.343 .